Keeley:

Raul bought a soft drink and a sandwich for 9.90. what is the price of each if the sandwich cost 3.5 times as much as the soft drink?

Keeley:

I am really stuck on this one could yall help?

generation2006:

ya just give me a min to type it

Keeley:

ok thank you

generation2006:

first create a ratio: the sandwich cost 3.5 more then the drink so it would be 3.5:1 once u did that your going to add the two sides of the ratio together to get the total part of the ratio 35+1=45 now that you know that total parts of the ratio you can solve fore one part by dividing the total cos, 9.90 by the total parts of the ratio, 4.5 9.90 divided by 4.5=2.20 since we know that each part of the ratio is 2.20 we multiply two sides of the ration by by 2.20 2.5*2.20=7.70 1*2.20=2.20 so then the answer would be B

generation2006:

hope it helped

generation2006:

and welcome 2 QC

jhonyy9:

so let the drink cost be an x and the sandwich cost be an y given again that total Raul bought a drink and a sandwich for $9,90 the sandwich cost 3,5 times as much as the drink so how you can write this using x and y ?

generation2006:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so let the drink cost be an x and the sandwich cost be an y given again that total Raul bought a drink and a sandwich for $9,90 the sandwich cost 3,5 times as much as the drink so how you can write this using x and y ? \(\color{#0cbb34}{\text{End of Quote}}\) hey did I get it right ur really good at math

jhonyy9:

ty

generation2006:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ty \(\color{#0cbb34}{\text{End of Quote}}\) What does ty mean again:")?

jhonyy9:

thank you

generation2006:

ur welcome but I don't know for what exactly

jhonyy9:

what you wrote above how i am in math

generation2006:

ur the expert i just do what i can